# LDM-ISP: Enhancing Neural ISP for Low Light with Latent Diffusion Models

### v3

@@ Abstract
Enhancing a low-light noisy RAW image into a well-exposed and clean sRGB image is a significant challenge for modern digital cameras. Prior approaches have difficulties in recovering fine-grained details and true colors of the scene under extremely low-light environments due to near-to-zero SNR. Meanwhile, diffusion models have shown significant progress towards general domain image generation. In this paper, we propose to leverage the pre-trained latent diffusion model to perform the neural ISP for enhancing extremely low-light images. Specifically, to tailor the pre-trained latent diffusion model to operate on the RAW domain, we train a set of lightweight taming modules to inject the RAW information into the diffusion denoising process via modulating the intermediate features of UNet. We further observe different roles of UNet denoising and decoder reconstruction in the latent diffusion model, which inspires us to decompose the low-light image enhancement task into latent-space low-frequency content generation and decoding-phase high-frequency detail maintenance. Through extensive experiments on representative datasets, we demonstrate our simple design not only achieves state-of-the-art performance in quantitative evaluations but also shows significant superiority in visual comparisons over strong baselines, which highlight the effectiveness of powerful generative priors for neural ISP under extremely low-light environments. Project page: https://csqiangwen.github.io/projects/ldm-isp/.
1 Introduction
@@
Enhancing low-light photos to well-exposed and clean RGB images not only requires reducing a significant amount of noise but also grapples with challenges related to detail and structure recovery, color correction, and careful joint optimization of many other operations. Thus, imaging in extremely low-light environments has been a long-standing challenging problem for modern digital cameras. Standard approaches such as increasing the ISO, extending exposure time, and enlarging the aperture during capturing, all have inevitable limitations and thus cannot solve the problem in practice. Learning-based methods, on the other hand, have been making surprising progress toward effective low-light imaging systems [2, 12, 38, 30, 28, 35, 1, 10, 41, 8, 11]. However, most of the existing approaches learn the low-light enhancement model from relatively limited paired data and the enhanced RGB images often suffer from detail deficiency and color deviation issues. In this paper, we leverage the powerful generative priors from a pre-trained latent diffusion model [21] to enhance the neural ISP for low-light imaging.
@@
Due to the significant practical value of low-light imaging, there have been many approaches to solve the task, which can be categorized into RGB-based enhancement [16, 15, 19, 37, 29, 33, 32] and RAW-based enhancement [2, 1, 30, 11, 28, 12, 38]. RGB data, however, contains much less signal than its unprocessed RAW counterparts, and thus the performance of RGB-based enhancement is limited, especially under extremely low-light conditions. To alleviate this issue, Chen et al. [2] propose to learn the entire ISP pipeline with paired low-light RAW and well-exposed sRGB data. Dong et al. [4] present to simulate the monochrome raw image from the colored raw image to compensate the photon number of the dark raw images. DNF [11] improves the neural ISP through decoupling the noise-to-clean and RAW-to-sRGB process, and utilizing the feature-level information of noise-to-clean as feedback to the RAW-to-sRGB process. Although the learned RAW-based low-light enhancement has been made effective progress, existing approaches still learn the enhancement process on limited pairs of RAW-sRGB data, and thus the resultant models can hardly perform well in extremely low-light cases.
Figure 1: An extremely low-light real-world image from SID Dataset [2] with a ×300 amplification ratio. The image processed by our proposed method exhibits clear improvement in structural content, showcasing finer details compared to previous works.($\dagger$ denotes the RAW-to-sRGB method and $\ast$ denotes the RAW-to-RAW method)
@@
2 Related Work
2.1 Learning-based Image Signal Processing
Deep learning has shown its advantages in learning the entire ISP process [23, 17, 36, 40, 31]. DeepISP [23] proposes one of the first CNN-based ISPs to replace the traditional human-tuned ISP operations, including denoising, demosaicking, color correction, white balance, and photo finishing steps. Instead of utilizing one neural network to simulate the entire ISP pipeline, CameraNet [17] presents a two-stage model for better RGB rendering performance. CycleISP [36] learns the bidirectional mapping of RGB-to-RAW and RAW-to-RGB to recover the RAW images from sRGB input. Recently, Xing et al. [31] proposes to utilize one single invertible neural network to learn an invertible ISP that can do both the RGB rendering and RAW reconstruction. However, most of the existing learning-based ISPs are designed for general photo-capturing conditions, and their performance under extremely low-light environments is unsatisfying. In this work, we aim to enhance the neural ISP for low light, with the abundant generative priors of pre-trained Stable Diffusion.
2.2 RAW-based Low-light Image Enhancement
@@
Compared to RAW-to-RAW methods, RAW-to-sRGB methods show a distinct advantage when the camera ISP is unknown or confidential. Chen et al[2] propose the first RAW-to-sRGB UNet-based method SID with a RAW dataset, which contains 5096 RAW data pairs for tackling the LLIE problem. However, this method exhibits severe color bias and a noticeable loss of details in extremely dark regions, primarily attributed to its simplistic single-stage network architecture. Zhu et al[41] propose a multi-stage method by fusing multi-exposure images to avoid the color bias issue. Xu et al[32] propose a multi-stage method based on a frequency-based decomposition to improve the detail enhancement. However, both methods suffer from error accumulation across stages [11]. Dong et al[4] propose a multi-stage method MCR that first predicts a monochrome raw image for information restoration and then performs LLIE. To address the challenge of the domain ambiguity issue in simultaneous denoising and ISP, Jin et al[11] propose a method DNF to tackle denoising and ISP tasks in separate stages. Although these multi-stage methods show promising LLIE performance, they still grapple with color correction and denoising in extremely challenging scenarios. The main reason for this shortcoming is that current training datasets [2, 30, 38, 4] contain limited scenes, which impedes the generalization ability of these learning-based methods.
In this work, instead of collecting a new dataset to improve the LLIE performance in extremely challenging scenarios, we propose to exploit generative priors from a powerful text-to-image latent diffusion model [21], which is pre-trained on a large-scale sRGB dataset, LAION-5B [22] containing 5.85 billion text-image pairs. With the help of generative priors learned from a large-scale open-domain dataset, the proposed method LDM-ISP not only presents satisfying denoising and ISP performance for low-light images, but also exhibits significant superiority in detail maintenance on extremely dark and noisy regions.
2.3 Diffusion Priors for Low-light Image Enhancement
With the rapid progress of diffusion models, the generative priors of diffusion models show promising potential for the LLIE task. Jiang et al[9] and Yi et al[34] propose diffusion model-based methods to perform sRGB-to-sRGB LLIE. Training a diffusion model from scratch on a small-scale LLIE dataset not only hampers the generalization ability of the method but also suffers from heavy computational costs. Meanwhile, Wang et al[28] propose to train a diffusion model from scratch for the RAW-to-RAW LLIE task. Based on the physical noise model, their diffusion model achieves low-light denoising in the RAW domain. Nevertheless, the challenges of the unknown ISP and unsatisfactory generalization problems persist in this method. Fei et al[5] propose to utilize the generative prior from a pre-trained diffusion model [3] for unified image restoration, including the LLIE task. By carefully designing the degradation model, their method regards the input low-light image as the guidance to control the generating process. Since the generated image is supervised by the input low-light image, the heavy noise from the input still remains in the final output. Although the above methods have promising potential for the LLIE task, they still suffer from noticeable artifacts or remaining noise.
In this work, instead of training or finetuning a diffusion model on an LLIE dataset, we insert a set of trainable taming modules into a frozen pre-trained latent diffusion model [21]. The generative prior of the latent diffusion model trained on a large-scale dataset helps the proposed method LDM-ISP achieve remarkable ISP enhancement performance for LLIE in diverse real-world challenging scenarios.
@@ 2.3 Diffusion Priors for Low-light Image Enhancement
Figure 2: The overview of our proposed LLIE method, LDM-ISP. The Bayer RAW image is processed by similar operations mentioned in [2]. A series of 2D discrete wavelet transforms (DWT) is applied to the processed image for capturing the low-frequency (LL) and high-frequency subbands (LH, HL, HH). The low-frequency subband (LL) serves to modulate the feature at each layer in the UNet. Specifically, each feature has its corresponding taming module, whose key part is an SFT layer, to map the sub-band into a pair of scale $\gamma$ and shift $\beta$ parameters. Similar to the low-frequency taming, the features of the decoder $\mathcal{D}$ are modulated by another set of taming modules, where the LL sub-band is mapped to the scale $\gamma$ and the concatenation of LH, HL, HH is mapped to the shift $\beta$. All parameters from the pre-trained Stable Diffusion are frozen and only taming modules are trainable.
3 Method
@@
An overview of the proposed method is shown in Fig. 2. Given a low-light noisy RAW image, we first process it to match the illumination to our target sRGB image: $X=(\text{RAW}-\text{Black Level})\times\gamma$, where $\gamma$ represents the user-defined amplification ratio. We aim at taming the latent diffusion model to perform the image signal processing pipeline (ISP) with restoration for over-dark areas and denoising for the processed image.
3.1 Generative Priors in the Latent Diffusion Model
@@
Figure 3: A text-to-image example from a well-known latent diffusion model, Stable Diffusion. The difference map between the latent representation and sRGB output indicates that the latent representation generated by the UNet shows essential structural content while the decoder $\mathcal{D}$ introduces abundant details but scarcely modifies the global structure for the final sRGB output. The latent representation is converted to sRGB visualization by the linear transformation [13].
@@
3.2 Low-frequency Content Generation
Figure 4: The low-frequency sub-band, extracted using 2D discrete wavelet transforms (DWT), reveals clearer structural information compared to the low-light input. On the other hand, merely resizing the input to align with the latent size of the UNet indicates that the information remains predominantly influenced by the noise. The result “Ours (Resize)” referring to merely resizing the input for taming the latent diffusion model shows over-smooth and poorly-structured results due to the negative impact of the noise.
Employing a series of DWT not only extracts the low-frequency sub-band (LL) but also allows us to efficiently align the input information with the latent size of the UNet. Compared to directly resizing the input RAW image, feeding the low-frequency sub-band (LL) to the UNet enables the model to concentrate on generating high-correlation structural content while reducing the negative impact of the noise, as shown in Fig. 4.
@@ 3.2 Low-frequency Content Generation
For the taming modules of the UNet, we employ a similar time-aware guidance strategy as discussed in [26]. However, instead of deriving the structural information by encoding the input image with the encoder $\mathcal{E}$ as in [26], we directly utilize the low-frequency sub-band (LL) to provide our structural information. This liberates us from the substantial computational cost of encoding, as shown in Table 1.
Table 1: A computational comparison between the encoder $\mathcal{E}$ adopted in [26] and our adopted DWT.
Input size
Time (s)
Params (M)
MACs (M)
Finetune for RAW
Encoder $\mathcal{E}$ $512\times 512$ 2.09 34.16 566,371
DWT $512\times 512$ 0.02 0 -
@@
3.3 High-frequency Detail Maintenance
This high-frequency “taming" aims to leverage the noise-free and fine-grained generative prior in the decoder to produce final sRGB images with enhanced details and devoided noise.
@@ 3.4 Training Objectives
Figure 5: The qualitative evaluations on SID-Sony [2] and ELD-Sony [30] datasets (Zoom-in for best view). The proposed method shows notable superiority in recovering structural information and enhancing details in extremely dark and noisy regions. ($\dagger$ denotes the RAW-to-sRGB method and $\ast$ denotes the RAW-to-RAW method)
Table 2: The quantitative comparison in extremely challenging scenarios from different datasets. The proposed method achieves state-of-the-art performance according to perceptual metrics, including LPIPS [39] and NIMA [25]. It indicates the sRGB results obtained from the proposed method exhibit significantly clearer global structure and more natural details compared to those from previous works. The best score is denoted as boldfaced.($\dagger$ denotes the RAW-to-sRGB method and $\ast$ denotes the RAW-to-RAW method)
Dataset
Amplification Ratio
Methods
LPIPS $\downarrow$
NIMA $\uparrow$
SSIM $\uparrow$
SID-Sony [2] $\times$300 ${}^{\dagger}$ [2] 0.3906 3.0481 0.8219
${}^{\ast}$ [30] 0.3876 3.0346 0.8271
${}^{\ast}$ [12] 0.4168 3.1831 0.8228
${}^{\dagger}$ [4] 0.4794 2.8077 0.7749
${}^{\dagger}$ [11] 0.4176 2.7646 0.8448
${}^{\dagger}$ 0.2749 3.5583 0.8113
ELD-Sony [30] $\times$200 ${}^{\dagger}$ [2] 0.2618 2.9780 0.7990
${}^{\ast}$ [30] 0.2585 3.1798 0.8063
${}^{\ast}$ [12] 0.2058 3.1718 0.8769
${}^{\dagger}$ [4] 0.3028 2.8887 0.7708
${}^{\dagger}$ [11] 0.2172 2.9123 0.8605
${}^{\dagger}$ 0.2035 3.4397 0.8906
LRD [38] -3EV ${}^{\ast}$ [38] 0.2198 3.1985 0.8784
${}^{\dagger}$ 0.1858 3.6180 0.8889
Figure 6: The qualitative evaluations on LRD [38] dataset (Zoom-in for best view).
4 Experiments
4.1 Implementation Details
The proposed method is evaluated on three widely-used datasets for the LLIE task, including SID-Sony [2], ELD-Sony [30], and LRD [38]. The train and test subsets follow the same strategy as previous works. Given the ground-truth RAW images from these datasets, we obtain their corresponding sRGB images through the open-source ISP pipeline, rawpy***https://pypi.org/project/rawpy/. The same ISP pipeline is also adopted for the outputs of our RAW-to-RAW baselines.
@@ 4.1 Implementation Details
4.2 Baselines
To demonstrate the superiority of the proposed LDM-ISP on enhancing ISP for low-light images, we first conduct a comparison with strong RAW-to-sRGB LLIE methods. To further highlight our denoising performance, we also compare the proposed method with RAW-to-RAW methods. The state-of-the-art methods compared with the proposed LDM-ISP are:
  • RAW-to-sRGB methods: SID [2], MCR [4] and DNF [11].
  • RAW-to-RAW methods: ELD [30], LED [12] and LRD [38].
4.3 Quantitative Evaluation
@@
According to Table  2, the proposed method outperforms other methods in terms of neural perceptual metrics, LPIPS [39] and NIMA [25]. To be specific, the proposed method achieves 0.1127 and 0.0023 improvements in LPIPS [39] than the previous sate-of-the-art methods on SID-Sony [2] and ELD-Sony [30] subsets, respectively. Meanwhile, there are 0.3752 and 0.2599 improvements in NIMA [25] on two subsets, respectively. It indicates the sRGB results obtained from the proposed method exhibit significantly more natural details compared to those from previous works. The structural quality assessment based on SSIM and LPIPS [39] indicates the proposed method has substantial advantages in enhancing low-frequency information, especially structural content. Note that the proposed method not only achieves state-of-the-art ISP performance compared to RAW-to-sRGB methods but also presents remarkably better denoising performance than RAW-to-RAW methods, which is a significant challenge to previous RAW-to-sRGB methods.
Figure 7: The visualization of the ablation study. Different taming strategies (resize and DWT) and taming targets (UNet and decoder $\mathcal{D}$) are combined in this study. Taming the UNet by resizing the input information tends to result in an ambiguous global structure. In contrast, the taming result using 2D discrete wavelet transforms (DWT) exhibits distinct structural content. Moreover, taming the decoder $\mathcal{D}$ results in improved color distribution and the enhancement of details.
4.4 Qualitative Evaluation
Fig. 5 and Fig. 6 show the visual comparison with other state-of-the-art methods on the SID-Sony [2], ELD-Sony [30], and LRD [38] datasets. Images captured in extremely low-light conditions require high amplification ratios ($\times$300 or -3 EV). These amplified images are plagued by acute noise and color distortion, making most details challenging to discern with the human eye. Compared with previous works, the proposed method shows notable superiority in recovering structural information and enhancing details in noisy regions. In the $\text{2}^{\text{nd}}$ column of Fig. 5, the results produced by previous works tend to be over-smooth due to the excessive denoising while the proposed method present sharp and clean leaves without color bias. In the $\text{5}^{\text{nd}}$ column of Fig. 5, previous works struggle to produce meaningful text information when portions of it are obscured by intense noise in the low-light image. The proposed method, which leverages the potent generative prior from pre-trained Stable Diffusion, exhibits a significantly improved reconstruction of the text.
Thanks to the noise-free generative prior in the decoder $\mathcal{D}$, the sRGB images produced by our proposed method consistently present clean content, irrespective of the degree of noise in low-light inputs.
4.5 Ablation Study
Table 3: The ablation study of the proposed method on ELD-Sony dataset [30]. Different taming strategies (resize and DWT) and taming targets (UNet and decoder $\mathcal{D}$) are combined in this study. The best score is denoted as boldfaced.
UNet
Decoder $\mathcal{D}$
SSIM $\uparrow$
LPIPS $\downarrow$

### v4

@@ Abstract
Enhancing a low-light noisy RAW image into a well-exposed and clean sRGB image is a significant challenge for modern digital cameras. Prior approaches have difficulties in recovering fine-grained details and true colors of the scene under extremely low-light environments due to near-to-zero SNR. Meanwhile, diffusion models have shown significant progress towards general domain image generation. In this paper, we propose to leverage the pre-trained latent diffusion model to perform the neural ISP for enhancing extremely low-light images. Specifically, to tailor the pre-trained latent diffusion model to operate on the RAW domain, we train a set of lightweight taming modules to inject the RAW information into the diffusion denoising process via modulating the intermediate features of UNet. We further observe different roles of UNet denoising and decoder reconstruction in the latent diffusion model, which inspires us to decompose the low-light image enhancement task into latent-space low-frequency content generation and decoding-phase high-frequency detail maintenance. Through extensive experiments on representative datasets, we demonstrate our simple design not only achieves state-of-the-art performance in quantitative evaluations but also shows significant superiority in visual comparisons over strong baselines, which highlight the effectiveness of powerful generative priors for neural ISP under extremely low-light environments.
I INTRODUCTION
@@
As the demand for robotic systems and surveillance technologies in low-light environments continues to rise, the importance of effective low-light image enhancement is becoming increasingly critical. Poor lighting conditions often lead to significant challenges in object detection, including increased noise and reduced visibility, which can severely impair performance. Enhancing low-light images is crucial for improving detection accuracy and reliability, enabling these systems to function effectively in diverse settings as shown in Fig. 1. Standard approaches such as increasing the ISO, extending exposure time, and enlarging the aperture during capturing, all have inevitable limitations and thus cannot solve the problem in practice. Learning-based methods, on the other hand, have been making surprising progress toward effective low-light imaging systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. However, most of the existing approaches learn the low-light enhancement model from relatively limited paired data and the enhanced RGB images often suffer from detail deficiency and color deviation issues.
Figure 1: The detection results obtained from YOLOv8 [12] demonstrate significant improvements from the proposed method. The proposed method yields clearer low-light enhancements, facilitating more accurate detection results.
@@
Due to the significant practical value of low-light imaging, there have been many approaches to solve the task, which can be categorized into RGB-based enhancement [13, 14, 15, 16, 17, 18, 19] and RAW-based enhancement [1, 7, 4, 11, 5, 2, 3]. RGB data, however, contains much less signal than its unprocessed RAW counterparts, and thus the performance of RGB-based enhancement is limited, especially under extremely low-light conditions. To alleviate this issue, Chen et al[1] propose to learn the entire ISP pipeline with paired low-light RAW and well-exposed sRGB data. Dong et al[20] present to simulate the monochrome raw image from the colored raw image to compensate the photon number of the dark raw images. DNF [11] improves the neural ISP through decoupling the noise-to-clean and RAW-to-sRGB process, and utilizing the feature-level information of noise-to-clean as feedback to the RAW-to-sRGB process. Although the learned RAW-based low-light enhancement has been made effective progress, existing approaches still learn the enhancement process on limited pairs of RAW-sRGB data, and thus the resultant models can hardly perform well in extremely low-light cases.
@@
II Related Work
II-A Learning-based Image Signal Processing
Deep learning has shown its advantages in learning the entire ISP process [24, 25, 26, 27, 28]. DeepISP [24] proposes one of the first CNN-based ISPs to replace the traditional human-tuned ISP operations, including denoising, demosaicking, color correction, white balance, and photo finishing steps. Instead of utilizing one neural network to simulate the entire ISP pipeline, CameraNet [25] presents a two-stage model for better RGB rendering performance. CycleISP [26] learns the bidirectional mapping of RGB-to-RAW and RAW-to-RGB to recover the RAW images from sRGB input. Recently, Xing et al[28] proposes to utilize one single invertible neural network to learn an invertible ISP that can do both the RGB rendering and RAW reconstruction. However, most of the existing learning-based ISPs are designed for general photo-capturing conditions, and their performance under extremely low-light environments is unsatisfying. In this work, we aim to enhance the neural ISP for low light, with the abundant generative priors of pre-trained Stable Diffusion.
II-B RAW-based Low-light Image Enhancement
@@
Compared to RAW-to-RAW methods, RAW-to-sRGB methods show a distinct advantage when the camera ISP is unknown or confidential. Chen et al[1] propose the first RAW-to-sRGB UNet-based method SID with a RAW dataset, which contains 5096 RAW data pairs for tackling the LLIE problem. Zhu et al[9] propose a multi-stage method by fusing multi-exposure images to avoid the color bias issue. Xu et al[19] propose a multi-stage method based on a frequency-based decomposition to improve the detail enhancement. However, both methods suffer from error accumulation across stages [11]. Dong et al[20] propose a multi-stage method MCR that first predicts a monochrome raw image for information restoration and then performs LLIE. To address the challenge of the domain ambiguity issue in simultaneous denoising and ISP, Jin et al[11] propose a method DNF to tackle denoising and ISP tasks in separate stages. Although these multi-stage methods show promising LLIE performance, they still grapple with color correction and denoising in extremely challenging scenarios. The main reason for this shortcoming is that current training datasets [1, 4, 3, 20] contain limited scenes, which impedes the generalization ability of these learning-based methods.
III Approach
@@
Figure 2: (a) The overview of our proposed LLIE method, LDM-ISP. The Bayer RAW image is processed by similar operations mentioned in [1]. A series of 2D discrete wavelet transforms (DWT) is applied to the processed image for capturing the low-frequency (LL) and high-frequency subbands (LH, HL, HH). The low-frequency subband (LL) serves to modulate the feature at each layer in the UNet. Specifically, each feature has its corresponding taming module, whose key part is an SFT layer, to map the sub-band into a pair of scale $\gamma$ and shift $\beta$ parameters. Similar to the low-frequency taming, the features of the decoder $\mathcal{D}$ are modulated by another set of taming modules, where the LL sub-band is mapped to the scale $\gamma$ and the concatenation of LH, HL, HH is mapped to the shift $\beta$. All parameters from the pre-trained Stable Diffusion are frozen and only taming modules are trainable. (b) The low-frequency sub-band, extracted using 2D discrete wavelet transforms (DWT), reveals clearer structural information compared to the low-light input.
@@
An overview of the proposed method is shown in Fig. 2 (a). Given a low-light noisy RAW image, we first process it to match the illumination to our target sRGB image. Then, we aim at taming the latent diffusion model to perform the image signal processing pipeline (ISP) with restoration for over-dark areas and denoising for the processed image.
@@
Figure 3: A text-to-image example from Stable Diffusion. The difference map between the latent representation and sRGB output indicates that the latent representation generated by the UNet shows essential structural content while the decoder $\mathcal{D}$ introduces abundant details but scarcely modifies the global structure for the final sRGB output. The latent representation is converted to sRGB visualization by the linear transformation [31].
III-A Generative Priors in the Latent Diffusion Model
@@
III-B Low-frequency Content Generation
Employing a series of DWT not only extracts the low-frequency sub-band (LL) but also allows us to efficiently align the input information with the latent size of the UNet. Compared to directly resizing the input RAW image, feeding the low-frequency sub-band (LL) to the UNet enables the model to concentrate on generating high-correlation structural content while reducing the negative impact of the noise, as shown in Fig. 2 (b).
@@ III-B Low-frequency Content Generation
For the taming modules of the UNet, we employ a similar time-aware guidance strategy as discussed in [34]. However, instead of deriving the structural information by encoding the input image with the encoder $\mathcal{E}$ as in [34], we directly utilize the low-frequency sub-band (LL) to provide our structural information. This liberates us from the substantial computational cost of encoding.
@@
III-C High-frequency Detail Maintenance
This high-frequency “taming” aims to leverage the noise-free and fine-grained generative prior in the decoder to produce final sRGB images with enhanced details and devoided noise.
@@ III-D Training Objectives
Figure 4: The qualitative evaluations on SID-Sony [1] and ELD-Sony [4] datasets (Zoom-in for best view). The proposed method shows notable superiority in recovering structural information and enhancing details in extremely dark and noisy regions. (The illumination of the low-light input is increased for visualization.)
IV Experiments
TABLE I: The quantitative comparison in extremely challenging scenarios from different datasets. The proposed method achieves state-of-the-art performance according to perceptual metrics, including LPIPS [36] and NIMA [37]. It indicates the sRGB results obtained from the proposed method exhibit significantly clearer global structure and more natural details compared to those from previous works. The best score is denoted as boldfaced.($\dagger$ denotes the RAW-to-sRGB method and $\ast$ denotes the RAW-to-RAW method)
Dataset Amplification Ratio Methods LPIPS $\downarrow$ NIMA $\uparrow$ SSIM $\uparrow$
SID-Sony [1] $\times$300 SID [1] 0.3906 3.0481 0.8219
ELD [4] 0.3876 3.0346 0.8271
LED [2] 0.4168 3.1831 0.8228
MCR [20] 0.5755 2.2367 0.7875
DNF [11] 0.5145 1.9610 0.8494
Ours 0.2749 3.5583 0.8113
ELD-Sony [4] $\times$200 SID [1] 0.2618 2.9780 0.7990
ELD [4] 0.2585 3.1798 0.8063
LED [2] 0.2058 3.1718 0.8769
MCR [20] 0.3120 2.2043 0.7798
DNF [11] 0.2255 2.0921 0.8683
Ours 0.2035 3.4397 0.8906
LRD [3] -3EV LRD [3] 0.2198 3.1985 0.8784
Ours 0.1858 3.6180 0.8889
IV-A Implementation Details
The proposed method is evaluated on three widely-used datasets for the LLIE task, including SID-Sony [1], ELD-Sony [4], and LRD [3]. The train and test subsets follow the same strategy as previous works. Given the ground-truth RAW images from these datasets, we obtain their corresponding sRGB images through the open-source ISP pipeline, rawpy [^1] https://pypi.org/project/rawpy/. The same ISP pipeline is also adopted for the outputs of our RAW-to-RAW baselines.
@@ IV-A Implementation Details
IV-B Baselines
To demonstrate the superiority of the proposed LDM-ISP on enhancing ISP for low-light images, we first conduct a comparison with strong RAW-to-sRGB LLIE methods. To further highlight our denoising performance, we also compare the proposed method with RAW-to-RAW methods. The state-of-the-art methods compared with the proposed LDM-ISP are: 1) RAW-to-sRGB methods including SID [1], MCR [20] and DNF [11]; 2) RAW-to-RAW methods including ELD [4], LED [2] and LRD [3].
Figure 5: The visualization of the ablation study. Different taming strategies (resize and DWT) and taming targets (UNet and decoder $\mathcal{D}$) are combined in this study. (The illumination of the low-light input is increased for visualization.)
IV-C Quantitative Evaluation
@@
According to Table  I, the proposed method outperforms other methods in terms of neural perceptual metrics, LPIPS [36] and NIMA [37]. To be specific, the proposed method achieves 0.1127 and 0.0023 improvements in LPIPS [36] than the previous sate-of-the-art methods on SID-Sony [1] and ELD-Sony [4] subsets, respectively. Meanwhile, there are 0.3752 and 0.2599 improvements in NIMA [37] on two subsets, respectively. It indicates the sRGB results obtained from the proposed method exhibit significantly more natural details compared to those from previous works. The structural quality assessment based on SSIM and LPIPS [36] indicates the proposed method has substantial advantages in enhancing low-frequency information, especially structural content. Note that the proposed method not only achieves state-of-the-art ISP performance compared to RAW-to-sRGB methods but also presents remarkably better denoising performance than RAW-to-RAW methods, which is a significant challenge to previous RAW-to-sRGB methods.
IV-D Qualitative Evaluation
Fig. 4 shows the visual comparison with other state-of-the-art methods on the SID-Sony [1], ELD-Sony [4] datasets. Images captured in extremely low-light conditions require high amplification ratios ($\times$300 or -3 EV). These amplified images are plagued by acute noise and color distortion, making most details challenging to discern with the human eye. Compared with previous works, the proposed method shows notable superiority in recovering structural information and enhancing details in noisy regions. In the $\text{1}^{\text{st}}$ row of Fig. 4, the results produced by previous works tend to be over-smooth due to the excessive denoising while the proposed method present sharp and clean leaves without color bias. In the $\text{3}^{\text{rd}}$ row of Fig. 4, previous works struggle to produce meaningful text information when portions of it are obscured by intense noise in the low-light image. The proposed method, which leverages the potent generative prior from pre-trained Stable Diffusion, exhibits a significantly improved reconstruction of the text.
IV-E Ablation Study
TABLE II: The ablation study of the proposed method on ELD-Sony dataset [4]. Different taming strategies (resize and DWT) and taming targets (UNet and decoder $\mathcal{D}$) are combined in this study. The best score is denoted as boldfaced.
UNet Decoder $\mathcal{D}$ SSIM $\uparrow$ LPIPS $\downarrow$